I have to find the coordinates of the vertical and the equation of the axis of symmetry for the parabola given by the q
Equation ;
Y= -1/4 x squared -1

Answer by lwsshak3(11628)

Find the coordinates of the vertex and the equation of the axis of symmetry for the parabola given.
Y= -1/4 x squared -1
y=-(1/4)x^2-1
4y=-x^2-4
x^2=-4y-4
x^2=-4(y+1)
This is an equation of a parabola that opens downwards
Its standard form: (x-h)^2=-4p(y-k), (h,k)=(x,y) coordinates of vertex
For given equation:x^2=-4(y+1)
vertex: (0,-1)
axis of symmetry: x=0 or y-axis